SOAR Issue Status

This page provides a brief summary of the status of significant issues with the SOAR Telescope and its associated instrumentation. The emphasis is on issues that are internal to SOAR. There are two tables - one comprising known issues, and the second second listing resolved issues in approximate reverse chronological order of resolution. The lists comprise only recurrent issues; responses to problems reported on individual night reports are not provided. In some cases there have been additional updates to "resolved" issues.

If a recurrent problem that you are aware of is not listed, please contact the instrument scientist, telescope scientist, or SOAR Director.

ACTIVE ISSUES

Issue

Description

Activity toward Resolution

Remote connection speed

Remote connections are slow, both for VNC and file downloads

Newer firewall now available with faster connections. NOTE - credentials subject to update as NOIRLab implements additional cybersecurity measures, but the faster speed is current available

SOI not in service

SOI is not available because a guider detector failed; can only be operated by “borrowing” the EM CCD from another guider (in this case, SOI)

Move CCD from optical Nasmyth now that new guider has been implemented there. Still pending.

Guider obsolescence

The facility guiders use obsolete hardware and do not support autofocus

The guider upgrade project is active, and installation of the new guider on the optical Nasmyth is complete.

Spartan IR Camera obsolescence

Spartan detectors are failing; overall performance is slow

The ISPI imager from Blanco is being upgraded and transferred to SOAR. Expected completion date mid-year 2023.

RESOLVED ISSUES (After Jun 2022)

Problem

Date Resolved

Description

Solution

Goodman internal flexure/jumps

Dec 2022 + Feb 2023

Spectra would shift or jump during longer observations

Visit from UNC repaired flexure compensation mechanism & trained local staff; updated GUI 

Goodman internal scattering

Dec 2022

Internal reflection from camera barrel at certain (low) grating angles compromised flat fields and spectra

Anodized camera barrel replaced with a barrel with a low-reflectivity coating

Elevation brake faults

Nov 2022

Mount control generated false brake fault alarms

Implemented relevant part of new mount control hardware

Elevation limit at 79 degrees

Nov 2022

Damage to encoder tape prevented telescope from operating above ~79 degrees elevation

New hardware (see previous) bypasses encoder that “sees” the damaged portion of the tape; a more permanent solution is pending.

Rotator brake faults

Aug 2022

Mount control generate false IR rotator brake faults

Bypassed alarm trigger; will be fully resolved with new mount control hardware

Updated on February 23, 2023, 8:16 am